# The uniqueness theorem for Kasparov theory

### v1

@@ 1. Preliminaries
Although we will introduce equivariant $KK$-theory in this section via the Cuntz–Thomsen picture [8, 9, 27, 42], we will implicitly assume that the reader has some existing passing familiarity with it; see for example [1]. We will assume that the reader is at least fluent in the language of $K$-theory for $\mathrm{C}^{*}$-algebras.
@@
Lemma 1.2 (see [31, Lemma 1.4] and its proof).
@@ 1.1. Some cocycle representation theory
We shall, from this point and throughout the article, always assume that $\beta:G\curvearrowright B$ denotes a strongly stable action on a $\sigma$-unital $\mathrm{C}^{*}$-algebra and that $\alpha:G\curvearrowright A$ is an action on a separable $\mathrm{C}^{*}$-algebra, unless specified otherwise.
@@
The last sentence of the following result and its justification follows the same line of reasoning as [8, Lemma 2.3] and [45, Lemma 6.1].
@@ 1.2. The Cuntz–Thomsen picture of equivariant $KK$-theory
such that the pointwise differences $\varphi-\psi$ and $\mathbbmss{u}-\mathbbmss{v}$ take values in $B\otimes\mathcal{K}$. [^4] In Thomsen’s article it was also assumed that the map $\mathbbmss{u}-\mathbbmss{v}$ is norm-continuous. This turns out to be redundant, see [41, Proposition 6.9]. A pair of the specific form $\big((\varphi,\mathbbmss{u}),(\varphi,\mathbbmss{u})\big)$ is called degenerate. Whenever $\beta$ is assumed to be strongly stable (this is our default assumption), we also allow $(B,\beta)$ in place of $(B\otimes\mathcal{K},\beta\otimes\operatorname{id}_{\mathcal{K}})$ appearing in the definition of an equivariant $(\alpha,\beta)$-Cuntz pair.
Notation 1.14.
@@
Definition 1.15 (see [42, Section 3] and [21, Section 2]).
@@
This Cuntz pair is independent of the choice of $t_{1},t_{2}$ up to homotopy; see [42, Lemma 3.4].
@@
The quotient $\mathbb{E}^{G}(\alpha,\beta)/{\sim_{h}}$ becomes an abelian group with Cuntz addition. The homotopy classes of cocycle pairs form a canonical subgroup $H_{\beta}$ independent of $\alpha$. It was proved by Thomsen in [42, Theorem 3.5] that the group quotient of $\mathbb{E}^{G}(\alpha,\beta)/{\sim_{h}}$ modulo $H_{\beta}$ is naturally isomorphic to $KK^{G}(\alpha,\beta)$ if one defines the latter via Kasparov’s original approach [31]. For an $(\alpha,\beta)$-Cuntz pair consisting of $(\varphi,\mathbbmss{u})$ and $(\psi,\mathbbmss{v})$, we denote its associated homotopy class by $(\varphi,\mathbbmss{u}),(\psi,\mathbbmss{v})]$. Under this identification, one has that the inclusion map $\mathbb{E}^{G}_{0}(\alpha,\beta)\subseteq\mathbb{E}^{G}(\alpha,\beta)$ also induces a natural isomorphism of abelian groups $\mathbb{E}^{G}_{0}(\alpha,\beta)/{\sim_{h}}\cong KK^{G}(\alpha,\beta)$. In other words, $KK^{G}(\alpha,\beta)$ may be defined as the abelian group of homotopy classes of anchored $(\alpha,\beta)$-Cuntz pairs.
@@ 2. $K_{1}$-injectivity of certain auxiliary $\mathrm{C}^{*}$-algebras
The following notation was already used in [21] and is inspired by notation appearing in the work of Dadarlat–Eilers in [12, Section 3]. We note that for $\mathrm{C}^{*}$-dynamics, a very similar type of auxiliary $\mathrm{C}^{*}$-algebras were considered by Thomsen in [45, Section 6] and their conceptual role was similar to how we will use the objects below in this article.
@@
Let $(\varphi,\mathbbmss{u}):(A,\alpha)\to(\mathcal{M}(B),\beta)$ be a cocycle representation. We then consider the two $\mathrm{C}^{*}$-algebras
@@
If $\varphi:A\to\mathcal{M}(B)$ is any *-homomorphism, then it can be viewed as a cocycle representation for actions of the trivial group, and we write
@@
If we assume that $(\varphi,\mathbbmss{u})$ contains itself at infinity and apply subsection 1.1, it follows that one has isomorphisms $\mathscr{D}^{w}_{(\varphi,\mathbbmss{u})}\cong\mathscr{D}^{w}_{(\varphi,\mathbbmss{u})^{\infty}}$ and $\mathscr{D}_{(\varphi,\mathbbmss{u})}\cong\mathscr{D}_{(\varphi,\mathbbmss{u})^{\infty}}$.
@@ 2. $\bullet$
The following is the main technical result of this section. A special case of it, stated after it, is the main ingredient to answer Introduction affirmatively, and its general form enables many of the results in the rest of the article.
@@
By section 2, we may assume without loss of generality that $(\varphi,\mathbbmss{u})$ is an infinite repeat. By subsection 1.1, we may assume that there exists a sequence of isometries $r_{n}\in(\mathcal{M}(B)\cap\varphi(A)^{\prime})^{\beta^{\mathbbmss{u}}}$ with $\mathbf{1}=\sum_{n=1}^{\infty}r_{n}r_{n}^{*}$. Depending on which case is proved, we write either $\mathscr{D}=\mathscr{D}_{(\varphi,\mathbbmss{u})}$ or $\mathscr{D}=\mathscr{D}_{(\varphi,\mathbbmss{u})}^{w}$, so that $\mathcal{E}=\mathscr{D}/\mathcal{J}$ in both cases. We shall assume that $\mathcal{E}$ is non-zero, otherwise there is nothing to prove.
We consider $p=P+\mathcal{J}$ for $P=\sum_{n=1}^{\infty}r_{2n}r_{2n}^{*}\in\mathscr{D}$, which defines a splitting projection in $\mathcal{E}$ with vanishing $K_{0}$-class (as $P$ is already such an element in $\mathscr{D}$). Let $q\in\mathcal{E}$ be another splitting projection with vanishing $K_{0}$-class. By section 2, the claim amounts to showing that $p$ and $q$ are homotopic. By section 2, $q$ is uniquely determined modulo unitary equivalence, thus there exists a unitary $w\in\mathcal{E}$ such that $w(\mathbf{1}-p)w^{*}=q$. Let $W\in\mathscr{D}$ be a contraction with $w=W+\mathcal{J}$. Set $Q=W(\mathbf{1}-P)W^{*}=W\big(\sum_{n=1}^{\infty}r_{2n-1}r_{2n-1}^{*}\big)W^{*}$, which then satisfies $q=Q+\mathcal{J}$.
@@
Consider the strict limits $\mathfrak{F}=\sum_{k=1}^{\infty}f_{k}$ and $\mathfrak{H}=\sum_{k=1}^{\infty}h_{k}$. Clearly $\mathfrak{F}\leq P$ and $\mathfrak{H}\leq Q$ and the above inequality implies $\|\mathfrak{F}\cdot\mathfrak{H}\|\leq\sum_{j,k=1}^{\infty}\|f_{k}h_{j}\|<1$.
@@
(e2.2) $\max_{g\in K_{k-2}}\|a_{k}-\beta_{g}^{\mathbbmss{u}}(a_{k})\|\leq 2^{-k},\quad\max_{a\in F_{k-2}}\|[a_{k},\varphi(a)]\|\leq 2^{-k}.$
@@
(e2.3) $\max_{g\in K_{k-2}}\|b_{k}-\beta_{g}^{\mathbbmss{u}}(b_{k})\|\leq 2^{-k},\quad\max_{a\in F_{k-2}}\|[b_{k},\varphi(a)]\|\leq 2^{-k}.$
@@
Condition (e2.2) implies for all $a\in\bigcup_{n}F_{n}$ that
@@
Likewise we can consider the strict limit $R_{2}^{\prime}=\sum_{k=1}^{\infty}r_{n_{k}}b_{k}$ and $R_{2}=WR_{2}^{\prime}$. As above, condition (e2.3) implies $R_{2}^{\prime}\in\mathscr{D}_{(\varphi,\mathbbmss{u})}\subseteq\mathscr{D}$ and thus $R_{2}\in\mathscr{D}$. Exactly as above it follows that $R_{2}^{\prime}$ is an isometry and hence
@@
To summarize, if we set $s_{j}=R_{j}+\mathcal{J}$ for $j=1,2$, then we have found two isometries $s_{1},s_{2}\in\mathcal{E}$ with
$s_{1}^{*}s_{1}^{*}\leq\mathfrak{F}+\mathcal{J}\leq P+\mathcal{J}=p,\quad s_{2}s_{2}^{*}\leq\mathfrak{H}+\mathcal{J}\leq Q+\mathcal{J}=q,$
@@
As a special case of this result, we obtain a positive solution to the $KK$-uniqueness problem for $\mathrm{C}^{*}$-algebras posed in recent work of Carrión et al; see also the thorough discussion following [39, Problem LXII]. Note that the special case $A=0$ recovers [19, Proposition 4.9], the special case $B=\mathcal{K}$ recovers [37, Lemma 3(2)], and [35, Theorem 2.5, 2.9], [36, Theorem 3.28] as well as [28, Corollary 4.2] are recovered as various other special cases. We note that the unitality condition written in brackets is an optional extra assumption one may add to the theorem in case $A$ is unital, so we obtain two versions of the statement below, with or without units. In particular we also obtain Introduction.
Corollary 2.9 (see [6, Question 5.17] and the surrounding discussion).
Let $A$ be a separable $\mathrm{C}^{*}$-algebra and $B$ a $\sigma$-unital stable $\mathrm{C}^{*}$-algebra. Let $\varphi:A\to\mathcal{M}(B)$ be a (unitally) absorbing representation and let $\bar{\varphi}:A\to\mathcal{Q}(B)=\mathcal{M}(B)/B$ be its induced homomorphism to the corona algebra. Then the relative commutant $\mathcal{Q}(B)\cap\bar{\varphi}(A)^{\prime}$ is $K_{1}$-injective. The same conclusion holds if $\varphi$ is instead assumed to be weakly nuclear and nuclearly (unitally) absorbing. In particular, Introduction has an affirmative answer.
@@
In the subsequent sections, we will generalize this consequence related to $KK$-uniqueness to the setting of generalized equivariant $KK$-groups with an approach that can be considered self-contained if one combines the methodology of this article with that of [21].
@@ 3. On a generalized equivariant $KK$-group construction
We apply the construction and argument of [21, Notation 3.7, Lemma 3.8]. This gives a cocycle representation $(\kappa,\mathbbmss{x})$ such that $(\varphi,\mathbbmss{u})\oplus(\kappa,\mathbbmss{x})$ and $(\psi,\mathbbmss{v})\oplus(\kappa,\mathbbmss{x})$ are operator homotopic. Upon a close inspection of how $(\kappa,\mathbbmss{x})$ is constructed there [^7] This is admittedly quite lengthy to check when reading, but we omit the details here as this just comes from following the relevant steps outlined in [21] verbatim., this cocycle representation arises as a Cuntz sum of $(\varphi,\mathbbmss{u})$ and a composition of cocycle representations that is of the same form as in section 3. To be more precise, if one starts with a representation $\theta_{0}:\mathcal{C}[0,1]\to\mathcal{M}(\mathcal{K}(\mathcal{H}))$ satisfying the conclusion of [21, Theorem 2.15], then one sets

### v2

@@ 1. Preliminaries
Although we will introduce equivariant $KK$-theory in this section via the Cuntz–Thomsen picture [8, 9, 27, 41], we will implicitly assume that the reader has some existing passing familiarity with it; see for example [1]. We will assume that the reader is at least fluent in the language of $K$-theory for $\mathrm{C}^{*}$-algebras. For the readers who are mainly interested in the results pertaining to $KK$-theory for $\mathrm{C}^{*}$-algebras (i.e., nonequivariant $KK$-theory), I wish to point out that one can choose to read the entire article by substituting $G=\{1\}$ and simply skip all the steps that involve the group actions.
@@
Lemma 1.2 (see [30, Lemma 1.4] and its proof).
@@ 1.1. Some cocycle representation theory
Convention 1.6.
We shall, from this point and throughout the article, always assume that $\beta:G\curvearrowright B$ denotes a strongly stable action on a $\sigma$-unital $\mathrm{C}^{*}$-algebra and that $\alpha:G\curvearrowright A$ is an action on a separable $\mathrm{C}^{*}$-algebra, unless specified otherwise. As we set out in section 1, $G$ is any second-countable locally compact group. We shall only repeat these assumptions sporadically and when stating the main results.
@@
The last sentence of the following result and its justification follows the same line of reasoning as [8, Lemma 2.3] and [44, Lemma 6.1].
@@ 1.2. The Cuntz–Thomsen picture of equivariant $KK$-theory
such that the pointwise differences $\varphi-\psi$ and $\mathbbmss{u}-\mathbbmss{v}$ take values in $B\otimes\mathcal{K}$. [^4] In Thomsen’s article it was also assumed that the map $\mathbbmss{u}-\mathbbmss{v}$ is norm-continuous. This turns out to be redundant, see [40, Proposition 6.9]. A pair of the specific form $\big((\varphi,\mathbbmss{u}),(\varphi,\mathbbmss{u})\big)$ is called degenerate. Whenever $\beta$ is assumed to be strongly stable (this is our default assumption), we also allow $(B,\beta)$ in place of $(B\otimes\mathcal{K},\beta\otimes\operatorname{id}_{\mathcal{K}})$ appearing in the definition of an equivariant $(\alpha,\beta)$-Cuntz pair.
Notation 1.15.
@@
Definition 1.16 (see [41, Section 3] and [21, Section 2]).
@@
This Cuntz pair is independent of the choice of $t_{1},t_{2}$ up to homotopy; see [41, Lemma 3.4].
@@
The quotient $\mathbb{E}^{G}(\alpha,\beta)/{\sim_{h}}$ becomes an abelian group with Cuntz addition. The homotopy classes of cocycle pairs form a canonical subgroup $H_{\beta}$ independent of $\alpha$. It was proved by Thomsen in [41, Theorem 3.5] that the group quotient of $\mathbb{E}^{G}(\alpha,\beta)/{\sim_{h}}$ modulo $H_{\beta}$ is naturally isomorphic to $KK^{G}(\alpha,\beta)$ if one defines the latter via Kasparov’s original approach [30]. For an $(\alpha,\beta)$-Cuntz pair consisting of $(\varphi,\mathbbmss{u})$ and $(\psi,\mathbbmss{v})$, we denote its associated homotopy class by $(\varphi,\mathbbmss{u}),(\psi,\mathbbmss{v})]$. Under this identification, one has that the inclusion map $\mathbb{E}^{G}_{0}(\alpha,\beta)\subseteq\mathbb{E}^{G}(\alpha,\beta)$ also induces a natural isomorphism of abelian groups $\mathbb{E}^{G}_{0}(\alpha,\beta)/{\sim_{h}}\cong KK^{G}(\alpha,\beta)$. In other words, $KK^{G}(\alpha,\beta)$ may be defined as the abelian group of homotopy classes of anchored $(\alpha,\beta)$-Cuntz pairs.
@@ 2. $K_{1}$-injectivity of certain auxiliary $\mathrm{C}^{*}$-algebras
The following notation was already used in [21] and is inspired by notation appearing in the work of Dadarlat–Eilers in [12, Section 3]. We note that for $\mathrm{C}^{*}$-dynamics, a very similar type of auxiliary $\mathrm{C}^{*}$-algebras was considered by Thomsen in [44, Section 6] and its conceptual role was similar to how we will use the objects below in this article.
@@
Let $\alpha:G\curvearrowright A$ and $\beta:G\curvearrowright B$ be actions on $\mathrm{C}^{*}$-algebras. Let $(\varphi,\mathbbmss{u}):(A,\alpha)\to(\mathcal{M}(B),\beta)$ be a cocycle representation. We then consider the two $\mathrm{C}^{*}$-algebras
@@
Note that these two objects coincide for $G=\{1\}$. Indeed, if $\varphi:A\to\mathcal{M}(B)$ is any *-homomorphism, then it can be viewed as a cocycle representation for actions of the trivial group, and we simply write
@@
For the rest of this section, we continue to follow subsection 1.1.
@@
If we assume that $(\varphi,\mathbbmss{u})$ contains itself at infinity and apply subsection 1.1, one obtains isomorphisms $\mathscr{D}^{w}_{(\varphi,\mathbbmss{u})}\cong\mathscr{D}^{w}_{(\varphi,\mathbbmss{u})^{\infty}}$ and $\mathscr{D}_{(\varphi,\mathbbmss{u})}\cong\mathscr{D}_{(\varphi,\mathbbmss{u})^{\infty}}$.
@@ 2. $\bullet$
The following is the main technical result of this section. A special case of it, stated after it, is the main ingredient to obtain Introduction, and its general form enables many of the results in the rest of the article.
@@
By section 2, we may assume without loss of generality that $(\varphi,\mathbbmss{u})$ is an infinite repeat. By subsection 1.1, we may assume that there exists a sequence of isometries $r_{n}\in(\mathcal{M}(B)\cap\varphi(A)^{\prime})^{\beta^{\mathbbmss{u}}}$ with $\mathbf{1}=\sum_{n=1}^{\infty}r_{n}r_{n}^{*}$. Depending on which case is proved, we write either $\mathscr{D}=\mathscr{D}_{(\varphi,\mathbbmss{u})}$ or $\mathscr{D}=\mathscr{D}_{(\varphi,\mathbbmss{u})}^{w}$, so that $\mathcal{E}=\mathscr{D}/\mathcal{J}$ in both cases. We assume that $\mathcal{E}$ is non-zero, else there is nothing to prove. We write $\pi_{\mathcal{J}}:\mathscr{D}\to\mathcal{E}$ for the quotient map.
We consider $p=\pi_{\mathcal{J}}(P)$ for $P=\sum_{n=1}^{\infty}r_{2n}r_{2n}^{*}\in(\mathcal{M}(B)\cap\varphi(A)^{\prime})^{\beta^{\mathbbmss{u}}}\subseteq\mathscr{D}$, which defines a splitting projection in $\mathcal{E}$ with vanishing $K_{0}$-class (as $P$ is already such an element in $\mathscr{D}$). Let $q\in\mathcal{E}$ be an arbitrary splitting projection with vanishing $K_{0}$-class. By section 2, the claim amounts to showing that $p$ and $q$ are homotopic. By section 2, $q$ is uniquely determined modulo unitary equivalence, thus there exists a unitary $w\in\mathcal{E}$ such that $w(\mathbf{1}-p)w^{*}=q$. Let $W\in\mathscr{D}$ be a contraction with $w=\pi_{\mathcal{J}}(W)$. Set $Q=W(\mathbf{1}-P)W^{*}=W\big(\sum_{n=1}^{\infty}r_{2n-1}r_{2n-1}^{*}\big)W^{*}$, which then satisfies $q=\pi_{\mathcal{J}}(Q)$.
@@
Consider the strict limits $\mathfrak{F}=\sum_{k=1}^{\infty}f_{k}$ and $\mathfrak{H}=\sum_{k=1}^{\infty}h_{k}$. Clearly $\mathfrak{F}\leq P$ and $\mathfrak{H}\leq Q$ and the above inequality implies
(e2.2) $\|\mathfrak{F}\cdot\mathfrak{H}\|\leq\sum_{j,k=1}^{\infty}\|f_{k}h_{j}\|<\sum_{j,k=1}^{\infty}2^{-(j+k)}=1.$
@@
(e2.3) $\max_{g\in K_{k-2}}\|a_{k}-\beta_{g}^{\mathbbmss{u}}(a_{k})\|\leq 2^{-k},\quad\max_{a\in F_{k-2}}\|[a_{k},\varphi(a)]\|\leq 2^{-k}.$
@@
(e2.4) $\max_{g\in K_{k-2}}\|b_{k}-\beta_{g}^{\mathbbmss{u}}(b_{k})\|\leq 2^{-k},\quad\max_{a\in F_{k-2}}\|[b_{k},\varphi(a)]\|\leq 2^{-k}.$
@@
Condition (e2.3) implies for all $a\in\bigcup_{n}F_{n}$ that
@@
Likewise we can consider the strict limit $R_{2}^{\prime}=\sum_{k=1}^{\infty}r_{n_{k}}b_{k}$ and $R_{2}=WR_{2}^{\prime}$. As above, condition (e2.4) implies $R_{2}^{\prime}\in\mathscr{D}_{(\varphi,\mathbbmss{u})}\subseteq\mathscr{D}$ and thus $R_{2}\in\mathscr{D}$. Exactly as above it follows that $R_{2}^{\prime}$ is an isometry and hence
@@
To summarize, if we set $s_{j}=\pi_{\mathcal{J}}(R_{j})$ for $j=1,2$, then we have found two isometries $s_{1},s_{2}\in\mathcal{E}$ with
$s_{1}^{*}s_{1}^{*}\leq\pi_{\mathcal{J}}(\mathfrak{F})\leq\pi_{\mathcal{J}}(P)=p,\quad s_{2}s_{2}^{*}\leq\pi_{\mathcal{J}}(\mathfrak{H})\leq\pi_{\mathcal{J}}(Q)=q,$
@@
We obtain the following as a consequence. Note that the special case $A=0$ recovers [19, Proposition 4.9], the special case $B=\mathcal{K}$ recovers [36, Lemma 3(2)], and [34, Theorem 2.5, 2.9], [35, Theorem 3.28] as well as [28, Corollary 4.2] are recovered as various other special cases. The unitality condition written in brackets is an optional extra assumption one may add to the theorem in case $A$ is unital, so we obtain two versions of the statement below, with or without units. In particular we also obtain Introduction.
Corollary 2.9.
Let $A$ be a separable $\mathrm{C}^{*}$-algebra and $B$ a $\sigma$-unital stable $\mathrm{C}^{*}$-algebra. Let $\varphi:A\to\mathcal{M}(B)$ be a (unitally) absorbing representation and let $\bar{\varphi}:A\to\mathcal{Q}(B)=\mathcal{M}(B)/B$ be its induced homomorphism to the corona algebra. Then the relative commutant $\mathcal{Q}(B)\cap\bar{\varphi}(A)^{\prime}$ is $K_{1}$-injective. The same conclusion holds if $\varphi$ is instead assumed to be weakly nuclear and nuclearly (unitally) absorbing.
@@
The above result implies a positive solution to the $KK$-uniqueness problem for $\mathrm{C}^{*}$-algebras posed in recent work of Carrión et al [6, Question 5.17]; see also [35, Theorem 2.6] or the discussion following [38, Problem LXII]. We record the resulting theorem here, whose separable version implies Introduction. Despite outsourcing the justification of the theorem to these references for the time being, we note that the subsequent sections will lead to a self-contained proof of Introduction, which is strictly more general than (i) below. Since the proof of [6, Theorem 5.15] makes it is clear that the $KL$-uniqueness theorem ((ii) below) can be reduced to (i), we omit a separate proof and shall not introduce or discuss $KL$-theory in the rest of the article.
Theorem 2.10.
Let $A$ be a separable $\mathrm{C}^{*}$-algebra and $B$ a $\sigma$-unital stable $\mathrm{C}^{*}$-algebra. Let $\varphi,\psi:A\to\mathcal{M}(B)$ be two absorbing *-homomorphisms that form a Cuntz pair.
  1. (i)
    If $\varphi,\psi]=0$ in $KK(A,B)$, then $\varphi$ and $\psi$ are strongly asymptotically unitarily equivalent.
  2. (ii)
    If $\varphi,\psi]=0$ in $KL(A,B)$, then there exists a sequence of unitaries $v_{n}\in\mathcal{U}(\mathbf{1}+B)$ such that
    $\varphi(a)=\lim_{n\to\infty}v_{n}\psi(a)v_{n}^{*},\quad a\in A.$
@@ 3. On a generalized equivariant $KK$-group construction
Throughout this section we continue following subsection 1.1.
@@
We apply the construction and argument as described in the statements of [21, Notation 3.7, Lemma 3.8]. This gives a cocycle representation $(\kappa,\mathbbmss{x})$ such that $(\varphi,\mathbbmss{u})\oplus(\kappa,\mathbbmss{x})$ and $(\psi,\mathbbmss{v})\oplus(\kappa,\mathbbmss{x})$ are operator homotopic. Upon a close inspection of how $(\kappa,\mathbbmss{x})$ is constructed there [^7] This is admittedly quite lengthy to check when reading, but we appeal to the construction described in [21, Notation 3.7] as-is. We omit the details here, as we would otherwise just copy content from this reference. Notably, the proof of “Lemma 3.8” therein is irrelevant for this discussion., this cocycle representation arises as a Cuntz sum of $(\varphi,\mathbbmss{u})$ and a composition of cocycle representations that is of the same form as in section 3. To be more precise, if one starts with a representation $\theta_{0}:\mathcal{C}[0,1]\to\mathcal{M}(\mathcal{K}(\mathcal{H}))$ satisfying the conclusion of [21, Theorem 2.15], then one sets